OpenStudy (mtalhahassan2):

Differentiate each of the following function: 2x^3-1/x^2

OpenStudy (mtalhahassan2):

F1(x)= 3(2x^2-1)(x^2)+(2x^3-1)(2x)

OpenStudy (mtalhahassan2):

So the derivative of the given function be like that right

OpenStudy (campbell_st):

well I'd start by writing both terms using index notation \[y = 2x^3 - x^{-2}\] now apply the rule for differentiation to both terms \[if~~ y = x^n ~~~y' = nx^{n - 1}\]

OpenStudy (mtalhahassan2):

Oh ok but you miss the -1

OpenStudy (mtalhahassan2):

2x^3-1

OpenStudy (campbell_st):

no I didn't I used index notation \[\frac{1}{x^2} = x^{-2}\]

OpenStudy (campbell_st):

so is the question 1. \[y = 2x^3 - \frac{1}{x^2}\] or 2. \[y = 2x^{3 - \frac{1}{x^2}}\]

OpenStudy (mtalhahassan2):

Both are wrong

OpenStudy (mtalhahassan2):

X^2 also a denominater of 2x^3

OpenStudy (mtalhahassan2):

@campbell_st

OpenStudy (campbell_st):

ok... so its \[\frac{2x^3 -1}{x^2}\]

OpenStudy (mtalhahassan2):

Yes

OpenStudy (mtalhahassan2):

F1(x)= 3(2x^2-1)(x^2)+(2x^3-1)(2x)

OpenStudy (campbell_st):

ok so there are a couple of ways to do this... 1. use the quitient rule... or 2. split it into 2 fractions... \[y = \frac{2x^3}{x^2} - \frac{1}{x^2}\] so simplifying and using index notation you get \[y = 2x - x^{-2}\] so the derivative is \[y' = 2 + \frac{2}{x^3}\]

OpenStudy (campbell_st):

using the method you ahve attempted, which looks like the quotient rule you forgot the denominator if \[y = \frac{f(x)}{g(x)} \] then \[y' = \frac{g(x) f'(x) - f(x) g'(x)}{(g(x))^2}\]

OpenStudy (mtalhahassan2):

No I guess that is product rule

OpenStudy (mtalhahassan2):

Y′=f(x)1h(x)+f(x)h(x)1

OpenStudy (mtalhahassan2):

@campbell_st

OpenStudy (campbell_st):

well if you are using the product rule you need to rewrite it using negative powers for the denominator \[\frac{2x^3 -1}{x^2} = (2x^3 -1)(x^{-2})\] which will make a big difference

OpenStudy (mtalhahassan2):

Oh ok but can I use any of them

OpenStudy (mtalhahassan2):

I mean the rules

OpenStudy (campbell_st):

you can use any of the rules, product rule, but careful with the power in the denominator quotient rule... normal diffierentiation rule after simplifying... they will all give the same answer

OpenStudy (mtalhahassan2):

Wait then what is ur answer

OpenStudy (mtalhahassan2):

F1(x)= 3(2x^2-1)(x)^-2+(2x^3-1)-2x

OpenStudy (mtalhahassan2):

F1(x)=6x^-2(x^2-1)-4x(x^3-1)

OpenStudy (mtalhahassan2):

@campbell_st

OpenStudy (anonymous):

You're making it harder than it is

OpenStudy (anonymous):

\[y=\frac{2x^3-1}{x^2} \iff y = (2x^3-1)(x^{-2})\]

OpenStudy (anonymous):

\[y'=2\cdot 3x^2 \cdot x^{-2} + (-2x^{-3})(2x^3-1)\]\[y'=2\cdot 3x^{2-2} -2x^{-3}(2x^3)+2x^{-3}\]\[y'=6x^{0}-(2\cdot 2)x^{0} +2x^{-3}\]\[y'=6-4+2x^{-3}\]\[y'= ...

OpenStudy (anonymous):

\[y' =....\] **

OpenStudy (triciaal):

do you have any questions?

OpenStudy (mtalhahassan2):

Can someone plz explain this question again to me??

OpenStudy (triciaal):

suggest you use parenthesis to make sure your question is clear

OpenStudy (triciaal):